40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10664     Entry time: Mon Nov 3 17:56:57 2014
Author: Koji 
Type: Update 
Category: LSC 
Subject: SRM calibration 

SRM Calibration

After the DRMI measurements on Friday, SRY cavity was locked in order to compare ITMY and SRM actuators.

SRY cavity was locked with AS55Q ->  SRM servo with gain of +10?
(My memory is fading. I tried +50 and noticed it was saturated at the limiter. So I thought it was 10)

Then the transfer functions between SRM->AS55Q TF and ITMY->AS55Q TF were measured.

The ratio between two transfer functions was obtained as seen in the second attachment.
The average at f<100Hz was 4.07 +/- 0.15. Therefore the calibration is ... as you can find below


SRM: http://nodus.ligo.caltech.edu:8080/40m/10664
SRM = (19.0 +/- 0.7) x 10 -9/ f2

PRM: http://nodus.ligo.caltech.edu:8080/40m/8255
PRM:  (19.6 +/- 0.3) x 10 -9 / f2 m/counts

BS/ITMs http://nodus.ligo.caltech.edu:8080/40m/8242
BS     = (20.7 +/- 0.1)    x 10 -9 / f2 m/counts
ITMX = (4.70 +/- 0.02)  x 10 -9/ f2
m/counts
ITMY = (4.66 +/- 0.02) x 10 -9/ f2
m/counts

Attachment 1: SRY_SRM_CALIB_RAW.pdf  1.224 MB  | Show | Show all
Attachment 2: SRY_SRM_CALIB.pdf  973 kB  | Show | Show all
ELOG V3.1.3-